OpenStudy (anonymous):

Evaluate 6^2 ÷ (3 + 9)

OpenStudy (anonymous):

36/12=3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

36/12 = 3

OpenStudy (anonymous):

(6*6) / (12) 36/12=3

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

that's what my calculator says.........